OpenStudy (anonymous):

If a and b are positive integers and their product is 3 times their sum, what is the value of (1/a) + (1/b)?

OpenStudy (anonymous):

\[ab = 3(a+b)\]

OpenStudy (anonymous):

I know how to set up the equation... I need help with the algebra LOL.

OpenStudy (anonymous):

ab=3(a+b) ==> a=3(a+b)/b ===> (1/a) + (1/b) = 1/[3(a+b)/b+1/b = b/3(a+b) + 1/b = [b^2+3a+3v]/[3ab+3b^2]

OpenStudy (anonymous):

Sorry but I can't read it... do you mind drawing it out? ><

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ab=3(a+b) ==> a=3(a+b)/b ===>\] \[(1/a) + (1/b) = 1/[3(a+b)/b+1/b\]\[ = b/3(a+b) + 1/b = [b^2+3a+3v]/[3ab+3b^2] \]

OpenStudy (anonymous):



OpenStudy (anonymous):

you may expand the bracjets also

OpenStudy (anonymous):

It's SAT math, so the answer has numerical digits :'( It's the free response.

OpenStudy (anonymous):

\[\frac{1}{a}+\frac{1}{b}=\frac{-3}{3b-ab}-\frac{3}{3a-ab}\]

OpenStudy (anonymous):

sin ab=3(a+b) divide ab on both side so (1/a+1/b)=1/3

OpenStudy (anonymous):

Thank you.